# Enhancing Robot Navigation Policies with Task-Specific Uncertainty Management

### v2

@@ 3.1. Task-Specific Uncertainty Map (TSUM)
To effectively integrate task-specific uncertainty considerations into robotic navigation, we introduce the concept of Task-Specific Uncertainty Maps (TSUMs). A TSUM is a function $U^{\tau}:L\rightarrow\mathbb{R}^{+}$ that assigns an acceptable level of state estimation uncertainty to each location $l\in L$ for a given navigation task $\tau$.
Formally, the TSUM $U^{\tau}(l)$ is defined as:
(1) $U^{\tau}(l)=\psi\left(\Phi^{\tau}(l),\mathcal{C}^{\tau}(l),\mathcal{E}(l)\right),$
where:
Task Relevance Function $\Phi^{\tau}(l)$:
This function quantifies the importance of location $l$ for the successful completion of task $\tau$. It measures how critical precise navigation and state estimation are at $l$ to fulfill the task objectives.
Task-Specific Constraints $\mathcal{C}^{\tau}(l)$:
These constraints represent conditions that must be satisfied at location $l$ for task $\tau$. They may include requirements such as avoiding restricted zones, operating within designated perimeters, or adhering to specific operational parameters unique to the task.
Environmental Factors $\mathcal{E}(l)$:
This component captures static properties of the environment at location $l$, such as the presence of obstacles, terrain difficulty, or regions known to be unsafe or infeasible for navigation.
The function $\psi$ combines these components into a scalar value $U^{\tau}(l)$, representing the acceptable level of uncertainty at location $l$ for task $\tau$. We define $\psi$ as a weighted sum:
(2) $U^{\tau}(l)=w_{\Phi}\Phi^{\tau}(l)+w_{\mathcal{C}}\mathcal{C}^{\tau}(l)+w_{\mathcal{E}}\mathcal{E}(l),$
where $w_{\Phi}$, $w_{\mathcal{C}}$, and $w_{\mathcal{E}}$ are weighting coefficients that determine the relative influence of each component.
We now present a method to generate TSUM from any given task description.
3.1.1. Implementation Details
To generate the Task-Specific Uncertainty Map (TSUM) $U^{\tau}(l)$, we combine task-specific semantics with spatial context.
Task Description Processing
Similar to the work in (Jie et al., 2017), we extract $m$ subtasks $\{t_{i}\}_{i=1}^{m}$ and $c$ constraints $\{c_{k}\}_{k=1}^{c}$ from the natural language task description $\tau$ using dependency parsing and Named Entity Recognition (NER).
Semantic and Spatial Embeddings
Subtasks and constraints are mapped into an $n$-dimensional semantic embedding space using the function $E_{c}:C\rightarrow\mathbb{R}^{n}$, implemented with a fine-tuned RoBERTa model (Liu, 2019) trained on navigation tasks:
(3) $E_{c}(t_{i}),\ E_{c}(c_{k})\in\mathbb{R}^{n}.$
Spatial locations $l\in L$ are characterized by their coordinates and environmental features. These features are processed by a neural network $f_{\text{spatial}}$ to produce spatial embeddings $E_{l}(l)\in\mathbb{R}^{n}$:
(4) $E_{l}(l)=f_{\text{spatial}}(\text{features}(l)).$
The network $f_{\text{spatial}}$ consists of two hidden layers with 128 and 64 units and ReLU activations.
Computing $\Phi^{\tau}(l)$ and $\mathcal{C}^{\tau}(l)$
Attention weights $\alpha_{i}$ and $\beta_{k}$ are assigned to subtasks and constraints using hierarchical attention networks (Yang et al., 2016):
(5) $\alpha_{i}=\frac{\exp(u_{i})}{\sum_{j=1}^{m}\exp(u_{j})},\quad u_{i}=v^{\top}\tanh(WE_{c}(t_{i})+b),$
(6) $\beta_{k}=\frac{\exp(u^{\prime}_{k})}{\sum_{k^{\prime}=1}^{c}\exp(u^{\prime}_{k^{\prime}})},\quad u^{\prime}_{k}=v^{\prime\top}\tanh(W^{\prime}E_{c}(c_{k})+b^{\prime}),$
where $W\in\mathbb{R}^{h\times n}$, $v\in\mathbb{R}^{h}$, $b\in\mathbb{R}^{h}$, $h$ is the hidden layer size and $W^{\prime}\in\mathbb{R}^{h^{\prime}\times n}$, $v^{\prime}\in\mathbb{R}^{h^{\prime}}$, $b^{\prime}\in\mathbb{R}^{h^{\prime}}$. The task relevance function $\Phi^{\tau}(l)$ and the task-specific constraints function $\mathcal{C}^{\tau}(l)$ are computed as:
(7) $\Phi^{\tau}(l)=\sum_{i=1}^{m}\alpha_{i}\cdot S_{c}\left(E_{c}(t_{i}),\ E_{l}(l)\right),$
(8) $\mathcal{C}^{\tau}(l)=\sum_{k=1}^{c}\beta_{k}\cdot S_{c}\left(E_{c}(c_{k}),\ E_{l}(l)\right),$
where $S_{c}(\cdot,\cdot)$ denotes the cosine similarity.
@@ 3.1.1. Implementation Details
Environmental factors are incorporated using a function $f_{\text{env}}$:
(9) $\mathcal{E}(l)=f_{\text{env}}(\text{env\_features}(l))=w_{\text{env}}^{\top}\text{env\_features}(l)+b_{\text{env}},$
@@
Aggregation into TSUM
The acceptable level of uncertainty at each location $l$ is determined by aggregating the computed functions using Equation (1).
3.1.2. Training the TSUM Generation Model
The embedding functions and attention networks are trained using a dataset of paired semantic concepts and spatial locations with known relationships [^1] The dataset was collected using an Autonomous Surface Vehicle (ASV) in a controlled environment. Details are provided in the appendix..
Triplet Loss for Embedding Alignment
To align the semantic and spatial embeddings, we employ a triplet loss function (Schroff et al., 2015):
(10) $L_{\text{triplet}}=\sum_{i=1}^{N}\left[\|E_{c}(t_{i})-E_{l}(l_{i}^{+})\|^{2}-\|E_{c}(t_{i})-E_{l}(l_{i}^{-})\|^{2}+\delta\right]_{+},$
where $N$ is the number of training samples, $l_{i}^{+}$ is a location relevant to subtask $t_{i}$, $l_{i}^{-}$ is an irrelevant location, $\delta$ is a margin parameter, and $[x]_{+}=\max(0,x)$.
This loss encourages the model to bring embeddings of related subtasks and locations closer while separating unrelated pairs.
Attention Network Training
The attention networks are trained by minimizing the Kullback-Leibler divergence (Kullback and Leibler, 1951) between the predicted attention distributions and the ground truth distributions:
(11) $L_{\text{attention}}=\text{KL}\left(\alpha\ \|\ \alpha^{\text{gt}}\right)+\text{KL}\left(\beta\ \|\ \beta^{\text{gt}}\right),$
where $\alpha^{\text{gt}}$ and $\beta^{\text{gt}}$ are the ground truth attention weights.
Total Loss Function
The total loss for training is:
(12) $L=L_{\text{triplet}}+\lambda L_{\text{attention}},$
with $\lambda$ being a hyperparameter balancing the two loss components.
Figure 3. Example task and its TSUM, where darker areas indicate a higher need for certainty and lighter areas signify a lower need for certainty.
@@ 3.1.2. Training the TSUM Generation Model
3.2. Conditioning Navigation Policies Using Task-Specific Uncertainty Maps
@@
(14) $\pi^{*}=\arg\max_{\pi}\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t},\tau)\,\Big{|}\,\tilde{s}_{0}\right],$
@@
In the next section, we show how a standard RL algorithm can be adapted using our framework to solve this optimization problem.
3.2.1. GUIDEd Soft Actor-Critic Algorithm
While any reinforcement learning algorithm can be used to learn the optimal policy $\pi^{*}(a|\tilde{s})$, we adopt the Soft Actor-Critic (SAC) algorithm (Haarnoja et al., 2018) due to its sample efficiency and robustness. Our adapted version, referred to as GUIDEd SAC, integrates TSUMs by conditioning the policy and value functions on the augmented state $\tilde{s}$.
@@ 3.2.1. GUIDEd Soft Actor-Critic Algorithm
(15) $J(\pi)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\left(R(s_{t},a_{t},\tau)+\alpha\mathcal{H}(\pi(\cdot|\tilde{s}_{t}))\right)\,\Big{|}\,\tilde{s}_{0}\right],$
@@
(16) $\mathcal{L}_{Q}(\phi_{i})=\mathbb{E}_{(\tilde{s}_{t},a_{t},r_{t},\tilde{s}_{t+1})\sim\mathcal{D}}\left[\left(Q_{\phi_{i}}(\tilde{s}_{t},a_{t})-y_{t}\right)^{2}\right],$
@@
(17) $y_{t}=r_{t}+\gamma\left(\min_{i=1,2}Q_{\bar{\phi}_{i}}(\tilde{s}_{t+1},a_{t+1})-\alpha\log\pi_{\theta}(a_{t+1}|\tilde{s}_{t+1})\right),$
@@
(18) $\mathcal{L}_{\pi}(\theta)=\mathbb{E}_{\tilde{s}_{t}\sim\mathcal{D}}\left[\mathbb{E}_{a_{t}\sim\pi_{\theta}(\cdot|\tilde{s}_{t})}\left[\alpha\log\pi_{\theta}(a_{t}|\tilde{s}_{t})-Q_{\phi}(\tilde{s}_{t},a_{t})\right]\right],$
@@
(19) $\mathcal{L}(\alpha)=\mathbb{E}_{a_{t}\sim\pi_{\theta}(\cdot|\tilde{s}_{t})}\left[-\alpha\left(\log\pi_{\theta}(a_{t}|\tilde{s}_{t})+\bar{\mathcal{H}}\right)\right],$
@@
4. Experiments
To evaluate the effectiveness of GUIDE, we conduct experiments in a real-world, real-time scenario using an autonomous surface vehicle (ASV). The ASV is tasked with navigating through a GPS-constrained environment, where precise localization comes at a cost. The agent operates with noisy position estimates by default and can request GPS readings when needed, reducing the uncertainty but at the cost of a higher penalty for each request.
The experimental setup is designed to mimic realistic operational constraints (Puthumanaillam et al., 2024; Zeng et al., 2018), highlighting the core challenge: balancing task success with effective uncertainty management. Excessive reliance on noisy localization risks failure in critical areas, while frequent use of high-precision GPS leads to increased mission costs.

### v3

@@ III-A Task-Specific Uncertainty Map (TSUM)
To effectively integrate task-specific uncertainty considerations into robotic navigation policies, we introduce the abstraction of Task-Specific Uncertainty Maps (TSUMs). A TSUM serves as an intermediate representation that bridges the gap between high-level task specifications and low-level uncertainty management decisions. Formally, a TSUM is a function $U^{\tau}:L\rightarrow\mathbb{R}^{+}$ that assigns an acceptable level of state estimation uncertainty to each location $l\in L$ for a given navigation task $\tau$.
The TSUM $U^{\tau}(l)$ is defined as:
$U^{\tau}(l)=w_{\Phi}\Phi^{\tau}(l)+w_{\mathcal{C}}\mathcal{C}^{\tau}(l)+w_{\mathcal{E}}\mathcal{E}(l),$ (1)
where $\Phi^{\tau}(l)$ quantifies the task-specific importance of location $l$, $\mathcal{C}^{\tau}(l)$ represents task constraints at $l$, and $\mathcal{E}(l)$ captures environmental factors. The weights $w_{\Phi}$, $w_{\mathcal{C}}$, and $w_{\mathcal{E}}$ control the relative influence of each component, allowing us to balance different aspects of the task and environment. These weights are determined through validation on a held-out dataset of annotated scenarios [^2] The dataset was collected using an Autonomous Surface Vehicle in a controlled environment. Details are provided in the supplementary. The dataset will be open-sourced after the review period.. By encoding the varying importance of uncertainty reduction across different regions of the environment, we provide a task-oriented abstraction of acceptable levels of state-estimation uncertainty that can be used to guide navigation policies.
III-A1 CLIP-Based TSUM Generation Pipeline
To generate TSUMs that effectively capture the relationship between visual environment features and task requirements, we propose a vision-language approach using the CLIP (Contrastive Language-Image Pre-training) model [50]. Our pipeline processes both visual and textual inputs to produce a spatially-varying uncertainty map. The following sections detail how we process environmental information and task specifications to compute each component of the TSUM, leveraging CLIP’s powerful cross-modal understanding capabilities.
Environment Representation
As established in Section III, we assume access to a visual representation of the operating environment through overhead imagery. To process this spatial information systematically, we discretize the continuous space $L$ into a grid of $N$ image patches $\{I_{1},I_{2},\ldots,I_{N}\}$, where each patch $I_{j}$ corresponds to a discrete spatial region $l_{j}\subset L$. The discretization parameters are chosen to balance computational efficiency with feature preservation.
Task Processing
Given a navigation task $\tau$ specified in natural language, we employ a hierarchical parsing approach to extract structured task representations. We utilize a dependency parser with a context-free grammar [55, 23] to decompose $\tau$ into atomic statements that capture both objectives and their associated constraints. These atomic statements are then organized into $m$ primary specifications $\{t_{i}\}_{i=1}^{m}$ and $c$ auxiliary specifications $\{c_{k}\}_{k=1}^{c}$, allowing for natural coupling between objectives and their execution requirements.
CLIP Encoding
We leverage CLIP’s vision-language architecture to project both spatial and semantic information into a shared representational space. The model consists of two parallel encoders:
$E_{\text{text}}:\mathcal{T}\rightarrow\mathbb{R}^{d},\quad E_{\text{image}}:\mathcal{I}\rightarrow\mathbb{R}^{d}$ (2)
where $\mathcal{T}$ and $\mathcal{I}$ denote the spaces of tokenized text and normalized image patches respectively, and $d=512$ matches CLIP’s native embedding dimension chosen to optimize the balance between representational capacity and computational efficiency. These encoders employ self-attention mechanisms and cross-modal transformers to generate context-aware embeddings $\mathbf{e}^{\text{(text)}}(t_{i}),\mathbf{e}^{\text{(text)}}(c_{k})\in\mathbb{R}^{d}$ for textual components and $\mathbf{e}^{\text{(image)}}(l_{j})\in\mathbb{R}^{d}$ for spatial regions.
Cross-Modal Alignment
The shared embedding space facilitates direct semantic comparison between task specifications and environmental features through cosine similarity [54]:
$\text{sim}(t,l)=\text{cosine}\big{(}\mathbf{e}^{\text{(text)}}(t),\;\mathbf{e}^{\text{(image)}}(l)\big{)}$
This metric quantifies the semantic relevance between spatial regions and task components, enabling the construction of spatially-varying relevance maps. While CLIP’s pre-training on diverse image-text pairs provides robust general-purpose understanding, domain-specific applications may benefit from additional fine-tuning. In our implementation, we fine-tune the last six transformer layers on a navigation dataset when deploying in maritime environments (details in Section III-A3).
The complete pipeline (illustrated in Fig. 3) processes environmental and task information in parallel streams, combining them through the CLIP embeddings. The following section details how these embeddings are used to compute task relevance and constraint components, which are then aggregated to generate the final TSUM representing acceptable uncertainty thresholds across the environment.
Figure 3: CLIP-based TSUM generation pipeline. The environment image and task description are processed through a fine-tuned CLIP model to obtain embeddings in a shared semantic space. These embeddings are used to compute attention-weighted task relevance and constraint components, which are combined with environmental factors to generate the final TSUM representing acceptable uncertainty levels across the environment.
III-A2 Computing Task Relevance and Constraints
Having established our CLIP-based pipeline, we now detail how we compute each component of the TSUM. Our approach leverages the rich semantic understanding captured by CLIP’s shared embedding space to quantify task relevance and constraints at each location.
Using the encoder functions defined in Equation (2), we obtain embeddings ($\mathbf{e}^{\text{(text)}}(t_{i}),\mathbf{e}^{\text{(text)}}(c_{k})\in\mathbb{R}^{d}$ for text and $\mathbf{e}^{\text{(image)}}(l_{j})\in\mathbb{R}^{d}$ for image patches) that capture the semantic features of both textual descriptions and visual characteristics in a shared $d$-dimensional space. This shared representation allows us to meaningfully compare how well visual features at each location align with the requirements expressed in the task description.
Task Relevance Function $\Phi^{\tau}(l)$
To compute the task relevance at each location, we first calculate similarity scores between each subtask and location
$\rho_{t_{i}}(l_{j})=\text{cosine}\big{(}\mathbf{e}^{\text{(text)}}(t_{i}),\;\mathbf{e}^{\text{(image)}}(l_{j})\big{)}.$ (3)
These raw similarities are converted to attention weights using a softmax operation
$\alpha_{i}(l_{j})=\frac{\exp\big{(}\rho_{t_{i}}(l_{j})\big{)}}{\sum_{i^{\prime}=1}^{m}\exp\big{(}\rho_{t_{i^{\prime}}}(l_{j})\big{)}}.$
The task relevance function $\Phi^{\tau}(l_{j})$ is then computed as a weighted sum
$\Phi^{\tau}(l_{j})=\sum_{i=1}^{m}\alpha_{i}(l_{j})\,\rho_{t_{i}}(l_{j}).$
This formulation creates a natural prioritization mechanism – locations that strongly match with important subtasks receive higher relevance scores. The attention weights $\alpha_{i}$ ensure that the most pertinent subtasks for each location dominate the relevance calculation, while the weighted sum combines evidence from all subtasks proportional to their importance.
Constraint Function $\mathcal{C}^{\tau}(l)$
We follow a similar procedure for computing constraint relevance. For each constraint $c_{k}$, we compute
$\rho_{c_{k}}(l_{j})=\text{cosine}\big{(}\mathbf{e}^{\text{(text)}}(c_{k}),\;\mathbf{e}^{\text{(image)}}(l_{j})\big{)}.$
Constraint attention weights are computed as
$\beta_{k}(l_{j})=\frac{\exp\big{(}\rho_{c_{k}}(l_{j})\big{)}}{\sum_{k^{\prime}=1}^{c}\exp\big{(}\rho_{c_{k^{\prime}}}(l_{j})\big{)}}.$
The final constraint function aggregates these weighted similarities
$\mathcal{C}^{\tau}(l_{j})=\sum_{k=1}^{c}\beta_{k}(l_{j})\,\rho_{c_{k}}(l_{j}).$
Higher values of $\mathcal{C}^{\tau}(l_{j})$ indicate locations where constraints are particularly relevant and uncertainty management is critical.
@@ III-A2 Computing Task Relevance and Constraints
$\mathcal{E}(l_{j})=w_{\text{env}}^{\top}\,\text{env\_features}(l_{j})+b_{\text{env}},$
where $w_{\text{env}}\in\mathbb{R}^{p}$ and $b_{\text{env}}\in\mathbb{R}$ are learned parameters optimized during training.
The final TSUM value at each location is computed by combining these components according to Equation (1).
III-A3 Implementation Details
We now present the practical details of our implementation, including dataset preparation, model fine-tuning, and training procedures.
Dataset
Our training dataset consists of 2,500 annotated overhead images collected from both real-world environments and high-fidelity simulations, covering a diverse range of navigation scenarios in a freshwater lake environment. Each image patch is discretized to 224 $\times$ 224 pixels at 0.5 meters per pixel resolution, chosen to align with CLIP’s architecture while preserving sufficient detail for navigation-relevant feature extraction. Each patch is annotated with:
$\mathcal{D}=\{(I_{j},l_{j},y_{j}^{t},y_{j}^{c},f_{j})\}_{j=1}^{N},$
where $I_{j}$ is the image patch, $l_{j}$ represents location coordinates, $y_{j}^{t}$ and $y_{j}^{c}$ are binary relevance labels for a predefined set of subtasks and constraints respectively, and $f_{j}$ contains environmental features including bathymetry, static obstacles, and known traffic patterns. The dataset is split into 2,000 images for training and 500 for validation.
CLIP Fine-tuning
We adopt a selective fine-tuning approach to adapt the pre-trained CLIP (ViT-B/32) model while preserving its general vision-language understanding. Specifically, we freeze the first six transformer layers to maintain CLIP’s foundational cross-modal understanding, while fine-tuning the remaining six layers to adapt to navigation-specific concepts. We optimize the following loss function:
$\mathcal{L}_{\text{CLIP}}=\mathcal{L}_{\text{contrast}}+\lambda_{\text{align}}\mathcal{L}_{\text{align}},$
where $\mathcal{L}_{\text{contrast}}$ is the standard CLIP contrastive loss
$\mathcal{L}_{\text{contrast}}=-\log\frac{\exp(\text{sim}(t_{i},l_{i}^{+})/\tau)}{\sum_{j}\exp(\text{sim}(t_{i},l_{j})/\tau)},$
and $\mathcal{L}_{\text{align}}$ is an additional alignment loss that encourages consistency between expert labels and embedding similarities
$\mathcal{L}_{\text{align}}=\sum_{j}\|y_{j}-\sigma(\text{sim}(t,l_{j})/\zeta)\|^{2}.$
Here, $\zeta$ is a temperature parameter set to 0.07, and $\sigma$ is the sigmoid function. We freeze the first six transformer layers of CLIP and fine-tune the remaining layers for 10 epochs using AdamW optimizer with a learning rate of 1e-5.
Training Procedure
The complete training process involves three stages:
  1. (i)
    We fine-tune CLIP using $\mathcal{L}_{\text{CLIP}}$ as described above.
  2. (ii)
    We train the environmental feature weights $w_{\text{env}}$ and $b_{\text{env}}$ by minimizing
    $\mathcal{L}_{\text{env}}=\sum_{j}\|\mathcal{E}(l_{j})-\hat{e}_{j}\|^{2},$
    where $\hat{e}_{j}$ represents ground truth annotations derived from auxiliary spatial data.
  3. (iii)
    Finally, we optimize the component weights $w_{\Phi}$, $w_{\mathcal{C}}$, and $w_{\mathcal{E}}$ using a validation set of TSUMs
    $\mathcal{L}_{\text{TSUM}}=\sum_{j}\|U^{\tau}(l_{j})-U^{\tau}_{\text{ref}}(l_{j})\|^{2}.$
The reference TSUMs ($U^{\tau}_{\text{ref}}$) were generated using a combination of automated and manual annotation processes. We utilized a high-fidelity Unity-based simulation environment to automatically generate initial uncertainty requirements based on task execution traces. These were then refined through a systematic review process to ensure alignment with navigation requirements. During training, we employ early stopping based on validation performance with a patience of 5 epochs. Data augmentation includes random rotations and flips of the image patches to improve robustness. The pipeline is implemented in PyTorch, with compute-intensive operations accelerated using TorchScript for optimal GPU utilization. Training is performed on NVIDIA A100 GPUs in a distributed setting using PyTorch’s DistributedDataParallel, with environment simulation and data generation parallelized across multiple nodes to maximize throughput.
@@ III-A3 Implementation Details
III-B Conditioning Navigation Policies Using Task-Specific Uncertainty Maps
@@
$\pi^{*}=\arg\max_{\pi}\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t},\tau)\,\Big{|}\,\tilde{s}_{0}\right],$
@@
Standard reinforcement learning algorithms can be employed to solve this optimization problem in the augmented state space. In the next section, we show how a standard RL algorithm can be adapted using our framework.
III-B1 GUIDEd Soft Actor-Critic Algorithm
While any reinforcement learning algorithm can be used to learn the optimal policy $\pi^{*}(a|\tilde{s})$, we adopt the Soft Actor-Critic (SAC) algorithm [17] due to its sample efficiency and robustness. Our adapted version, referred to as GUIDEd SAC, integrates TSUMs by conditioning the policy and value functions on the augmented state $\tilde{s}$.
@@ III-B1 GUIDEd Soft Actor-Critic Algorithm
$J(\pi)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\left(R(s_{t},a_{t},\tau)+\alpha\mathcal{H}(\pi(\cdot|\tilde{s}_{t}))\right)\,\Big{|}\,\tilde{s}_{0}\right],$
@@
$\mathcal{L}_{Q}(\phi_{i})=\mathbb{E}_{(\tilde{s}_{t},a_{t},r_{t},\tilde{s}_{t+1})\sim\mathcal{D}}\left[\left(Q_{\phi_{i}}(\tilde{s}_{t},a_{t})-y_{t}\right)^{2}\right],$ (4)
@@
$y_{t}=r_{t}+\gamma\left(\min_{i=1,2}Q_{\bar{\phi}_{i}}(\tilde{s}_{t+1},a_{t+1})-\alpha\log\pi_{\theta}(a_{t+1}|\tilde{s}_{t+1})\right),$ (5)
@@
$\mathcal{L}_{\pi}(\theta)=\mathbb{E}_{\tilde{s}_{t}\sim\mathcal{D}}\left[\mathbb{E}_{a_{t}\sim\pi_{\theta}(\cdot|\tilde{s}_{t})}\left[\alpha\log\pi_{\theta}(a_{t}|\tilde{s}_{t})-Q_{\phi}(\tilde{s}_{t},a_{t})\right]\right],$ (6)
@@
$\mathcal{L}(\alpha)=\mathbb{E}_{a_{t}\sim\pi_{\theta}(\cdot|\tilde{s}_{t})}\left[-\alpha\left(\log\pi_{\theta}(a_{t}|\tilde{s}_{t})+\bar{\mathcal{H}}\right)\right],$ (7)
@@
IV Experiments
Motivating scenario: Consider a stealth reconnaissance mission where an autonomous vehicle must navigate through previously unexplored waters. The vehicle can either rely on passive, noisy sensors or risk detection by activating high-precision GPS. Each GPS usage increases the chance of compromising the mission, yet completing the assigned tasks requires maintaining sufficient localization accuracy in critical regions.
To evaluate GUIDEd agents in such challenging scenarios, we conduct experiments using an autonomous surface vehicle (ASV) in real-world, previously unseen environments. The ASV is tasked with executing a set of navigation tasks that it has never encountered during training. Operating in a GPS-constrained setting, the ASV must strategically manage its localization resources: by default, it relies on passive sensors providing noisy position estimates, but can request precise GPS fixes at the cost of increased mission risk. The experimental setup is designed to capture realistic operational constraints [49, 72] where robots must balance competing objectives. Relying solely on noisy localization can lead to task failure in regions requiring precision, while excessive use of GPS not only increases mission costs but could compromise stealth objectives.